OpenStudy (anonymous):

Use the quadratic formula to solve: x^2-28=0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (paki):

@hrassistant hey welcome to open study... please respond here, so that we may guide you...